8=sqrt(x)sqrt(x)sqrt(x)sqrt(x...), find the value of x.
Answer by RAY100(1637)

8 =sqrtx sqrtx sqrtx sqrtx
.
8 = (x^1/2)^4 = x^(4/2) = x^2
.
sqrt8 = 2.828 = x
.
checking
.
8 = (sqrt 2.828)^4 = 8,,, ok